OpenStudy (anonymous):

solve for x by completing the square 1. x^2 - mx - 6m^2 = 0 2. kx^2 + px + q = 0

OpenStudy (anonymous):

\[\begin{align*}x^2-mx-6m^2&=0\\ x^2-mx+\frac{m^2}{4}-\frac{m^2}{4}&=6m^2\\ \left(x-\frac{m}{2}\right)^2&=6m^2+\frac{m^2}{4}\\ \left(x-\frac{m}{2}\right)^2&=\frac{23m^2}{4}\\ x-\frac{m}{2}&=\pm\sqrt{\frac{23m^2}{4}}\\ x&=\frac{m}{2}\pm\frac{\sqrt{23m^2}}{2}\\ \end{align*}\]